OpenStudy (anonymous):

Find the derivative of f(t) = sqrt((2t+1)/(t^2+2))

OpenStudy (anonymous):

you're going to have to either use product by bringing your denominator up or use quotient... which do you want to

OpenStudy (anonymous):

... not getting answer must be afk

myininaya (myininaya):

(y=f(t)) i like to find y' here by taking natural log of both sides

OpenStudy (anonymous):

interesting lets see where you are getting at this lol

myininaya (myininaya):

\[\ln(y)=\ln(\sqrt{\frac{2t+1}{t^2+2}})\] \[\ln(y)=\frac{1}{2} [\ln|2t+1|-\ln(t^2+2)]\] now find y' by doing first taking derivative of both sides \[\frac{y'}{y}=\frac{1}{2}[\frac{2}{2t+1}-\frac{2t}{t^2+2}]\]

myininaya (myininaya):

now simply multiply y on both sides

myininaya (myininaya):

and replace y with \[\sqrt{\frac{2t+1}{t^2+2}} \]

myininaya (myininaya):

and done

OpenStudy (anonymous):

-(2/((t²+2)²))(t²+t-2) is the answer

myininaya (myininaya):

if you want you can use algebra to show that if that is indeed the answer

myininaya (myininaya):

take my expression and make it into that expression by using algebra manipulation

OpenStudy (anonymous):

i didnt see sqrt. -(1/(√(2t+1)(t²+2)²))(3t²+2t-2)

OpenStudy (anonymous):

this must be correct ans

OpenStudy (anonymous):

thanks guys this problem is a pain

OpenStudy (anonymous):

really this problem shouldn't be that hard if you just move the bottom up by giving it a negative exponent and do the product rule basically with something like this instead of the quotient rule which is just the product rule written another way, i go through these steps 1)Take the denominator and make it a numerator with negative exponent 2)Do product rule 3)After product rule, move the negative exponents back to the denominator 4)Usually after this step you can easily make a common denominator, do that 5)Add like terms and simplify the numerator 6)Profit

OpenStudy (anonymous):

your way to approach this problem is very good, thanks for the help

myininaya (myininaya):

i still like my way much more :)